A company expects to borrow $1 million in money markets in three months’ time. The current interest rates are 2.40% and expected to rise. The company uses the futures markets to sell ten bank bill futures at 97.50. Calculate the value of ten bank bill futures contract when one futures contract has a face value of $1,000,000.

Answer: $993873.38

How to work it out please?
Answer #1

A Bank Bill futures contract is trading at 97.50,

Yield = 100-95.50 = 2.5%

Value of bank bill ,

P = Face vale * 365 365+ yield-90

P = 1000000*365/(365+(2.5*90/100)) = $ 993,873.38.